OpenStudy (anonymous):

FAN AND METAL Find an equation in standard form for the hyperbola with vertices at (0, ±3) and foci at (0, ±7).

OpenStudy (amistre64):

any ideas how to proceed?

OpenStudy (anonymous):

no idk how to do it can you explain

OpenStudy (amistre64):

your material should give you a standard general setup for the equation that we can start to fill in. we should start with that.

OpenStudy (amistre64):

we will want to find a center point, and distances from that center to a vertex, a focus, and the rest is pythag thrm

OpenStudy (anonymous):

so how would i start

OpenStudy (amistre64):

your material should give you a standard general setup for the equation that we can start to fill in. we should start with that.

OpenStudy (amistre64):

if you cant find it in your lesson material, the internet has a plethera of them laying about

OpenStudy (anonymous):

(x-h)^2/a^2 -(y-k)^2/b2

OpenStudy (anonymous):

thats what it gives me

OpenStudy (amistre64):

that one will do fine. now the point (h,k) defines the center point. that will be the midpoint between vertexes, or the midpoint between focuses

OpenStudy (amistre64):

and for a little correction, we need the equation to equal 1 (x-h)^2/a^2 -(y-k)^2/b2 = 1 can you find the center point for us?

OpenStudy (anonymous):

how do i do that

OpenStudy (amistre64):

the center point is an average of the extremes: lets use the verts: (0,3) and (0,-3) now if its not obvious that (0,0) is the center between them we can go about it as: 0+0 3-3 ---- , --- = (0,0) 2 2

OpenStudy (amistre64):

adding the component parts, and dividing by 2

OpenStudy (anonymous):

okay so the center is (0,0)

OpenStudy (amistre64):

yes, which tells us (h,k) = (0,0), so lets use that in the setup x^2/a^2 - y^2/b^2 = 1 now a hyperbola, is not necessarily x-y, but maybe y-x so we need to determine the proper sign placement we are told that when x=0, y=7 by the vertex, agreed? using that we get: 0/a^2 - 7^2/b^2 = 1; now -7^2/b^2 can never equal +1 can it? so we must change the order of the setup to: y^2/b^2 - x^2/a^2 = 1 does this make sense?

OpenStudy (amistre64):

i should have used the verts on that, but at my age i tend to get things mixed about ....

OpenStudy (anonymous):

not really

OpenStudy (amistre64):

what about it is not making sense?

OpenStudy (amistre64):

does it make sense up to a point? or is it all just muddled?

OpenStudy (amistre64):

do we agree that h=0, and k=0?

OpenStudy (anonymous):

yea i get the center

OpenStudy (amistre64):

\[\pm\frac{(x-0)^2}{a^2}\mp\frac{(y-k)^2}{b^2}=1\] \[\pm\frac{x^2}{a^2}\mp\frac{y^2}{b^2}=1\] we are given that the point (0,3) has to be a solution to the equation, we can use that to determine 2 things, sign usage and one of the values for a^2 or b^2 \[\mp\frac{3^2}{b^2}=1\]since a negative is never equal to a postive, that means the sign is a +, and the only way this can be equal to 1 is if b^2 =3^2 \[-\frac{x^2}{a^2}+\frac{y^2}{3^2}=1\]or written a more standard way \[\frac{y^2}{3^2}-\frac{x^2}{a^2}=1\] and we have only one more thing to determine to finish it up

OpenStudy (amistre64):

in a hyperbola, the pythagorean thrm comes into play; if we call f the distance from the center to a focus point: then a^2 + b^2 = f^2 and we can solve for a^2 a^2 = f^2 - b^2 a^2 = 7^2 - 3^2